OpenStudy (anonymous):

what is the answer to 80-6x=8

OpenStudy (anonymous):

\[80-6x=8\] First subtract 80 from both sides; \[80-6x-80=8-80\] Which leaves; \[-6x=-72\] Next, divide both sides by -6; \[\frac{ -6x }{ -6 }=\frac{ -72 }{ -6 }\] This lets you get x alone. \[x=12\]